# A faithful action of $\overline{\mathbb{Q}}/\mathbb{Q}$) on Zariski multiplets

### v1

@@ 1. Introduction
The notion of Zariski pair was introduced by E. Artal in his seminal paper [A94]. This put a new focus on an important question for plane complex projective curves:
Do the topological types of the singularities of a plane curve $C$ and the degree determine the homeomorphism type of the pair $\mathbb{P}^{2},C$?
@@
In our previews work [LP20] we used the fundamental group of the complement together with the conjugacy class of meridians to the curve, which is well-defined for irreducible curves, in the second step.
@@
In particular we can derive from that set-up the existence of arbitrarily large arithmetic Zariski multiplets of plane curves.
Theorem 1.3.
There exist arithmetic Zariski multiplets of equisingular plane irreducible curves with at most nodal and ordinary cuspidal singularities of arbitrary large finite cardinality.
While this existence result does not come with estimates for the numerical invariants, we are able to make use of specific examples of large Galois orbits of surfaces for the final result.
@@
$\deg(p_{i})=60p_{i}\left(\frac{p_{i}-1}{6}-1\right)\left(\frac{p_{i}+1}{4}-1\right).$
@@
In the final section we first demonstrate that the associated generic $m$-canonical branch curves of surfaces isogenous to a product with different fundamental group form a Zariski pair (see Theorem 3.7) and, using, in particular cases, the functor $F_{\sigma}$, even arithmetic Zariski pairs. Finally, we provide detailed proofs of the three main theorems stated in the introduction.
@@ 2. Preliminaries
$\begin{split}F_{\sigma}\colon\underline{\mathbf{\mathbb{C}-var}}\longrightarrow&\underline{\mathbf{\mathbb{C}-var}}\\ X\mapsto&X^{\sigma}\\ \left(f\colon X\longrightarrow Y\right)\mapsto&\left(f^{\sigma}\colon X^{\sigma}\longrightarrow Y^{\sigma}\right)\end{split}$
@@
and the family $\mathcal{X}^{\sigma}\longrightarrow B^{\sigma}$ connects the surfaces $S_{1}^{\sigma}$ and $S_{2}^{\sigma}$. More precisely, choose the canonical model $X$ of a surface of general type $S$ and apply the field automorphism $\sigma\in{\mathrm{Aut}}(\mathbb{C})$ to a point of the Hilbert scheme corresponding to the $m$-canonical image of $S$ We obtain a surface $X^{\sigma}$, and whose minimal model is $S^{\sigma}$. Since the Hilbert scheme corresponding to $m$-pluricanonical embedded surfaces is defined over $\mathbb{Z}$, it is invariant under the functor $F_{\sigma}$ so its connected components are defined over $\overline{\mathbb{Q}}$. We conclude that the elements in the kernel of $\varphi\colon{\mathrm{Aut}}(\mathbb{C})\longrightarrow{\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}$ act trivially on connected components of the moduli space of surfaces which is in bijection with the set of connected components of the Hilbert subscheme just defined. For more details see [BCG15, Proposition 5.7].
We want to do something similar for parameter spaces of plane curves. More precisely, we can consider equisingularity strata ( they replace the Hilbert scheme ) of plane curves with a fixed combinatorial type. Like the Hilbert scheme, such strata are in general neither irreducible nor connected.
The combinatorial type of plane curves has been defined in [ACT10] in full generality. We only consider the relevant restriction to irreducible curves, that is, the number of singularities of each topological type and the degree.
@@
Given a combinatorial type of degree $d$ plane curves the corresponding stratum in the parameter space of degree $d$ plane curve is given by the locus of all curves with that combinatorial type.
The easiest types of singularities are ordinary nodes and cusps characterized by their Milnor numbers for examples.
@@
Let $B_{1}$ and $B_{2}$ two equisingular plane curves we call them rigidly isotopic if there is a family $\mathcal{B}\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}B$ of equisingular curves over a smooth base $B$ such that there exist $b_{1}$ and $b_{2}$ in $B$ such that $\pi^{-1}(b_{i})=B_{i}$ for $i=1,2$.
Hence connected components of strata are the same as rigid isotopy classes. In a similar way as for surfaces we use the diagram (1) and get:
@@
We want to link more tight the action of ${\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}$ on the connected components of moduli space of surfaces with the action on the connected components of equisingular strata of plane curves with only nodes and cusps as singularities. To do so we have to recall the general theory of branched curves, this is the same strategy we used in our previous paper [LP20].
@@
${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{m}}$$\scriptstyle{p}$${\mathbb{P}^{P_{m}-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{P}^{2}}$
@@
We are now able to associate to a surface of general type a plane curve with only nodes and cusps as singularities. Among all the surfaces of general type there are some for which is possible to control accurately the ${\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}$ action on the connected components of their moduli space. These surfaces are the so called surfaces isogenous to a product (or SIPs for short). They were introduced by Catanese in [Cat00] and the action of ${\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}$ on their moduli space was investigated in [BCG15].
@@
(4) $1\longrightarrow\pi_{1}(C_{1})\times\pi_{1}(C_{2})\longrightarrow\pi_{1}(S)\longrightarrow G\longrightarrow 1$
@@
[…]
@@ 4. Proof of the main theorems
By Theorem 3.7 $(\mathbb{P}^{2},B)$, $(\mathbb{P}^{2},B^{\sigma})$ are Zariski pair. We only have to prove that the pair is arithmetic. By the proof of [BCG15, Theorem 1.3] the field of moduli is a number field and so is a field of definition therefore we know that the $m$-canonical model $X$ of $S$ and $X^{\sigma}$ for $S^{\sigma}$ in $\mathbb{P}^{P_{m}-1}$ is defined over $\overline{\mathbb{Q}}$. By [CiF11, Theorem 1.1] we can map $S$ to $\mathbb{P}^{2}$ by a generic projection $\pi$ with branch curve $B$ that is irreducible with at most nodes and ordinary cusps as singularities and again defined over a number field. In particular, this allows us to apply the functor $F_{\sigma}$ defined in 2.2 and to obtain the following diagram.
[…]
@@
We are now in the position to prove Theorems 1.2, 1.3, 1.4, that we recall for reader’s convenience.
@@
Theorem 4.3.
There exist arithmetic Zariski multiplets of equisingular plane irreducible curves with at most nodal and ordinary cuspidal singularities of arbitrary large finite cardinality.
Proof.
Fix a prime number $p$. There is $\sigma\in{\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}$ of order $p$ which extends to an automorphism $\tilde{\sigma}$ of the field $\mathbb{C}$. By [BCG15] there exist SIP surfaces $X$ and $X^{\sigma}$ that have non isomorphic topological fundamental group. By our standard strategy we embed $X$ by the $3-$canonical map into a projective space in such a way that the image is defined over $\overline{\mathbb{Q}}$. Indeed, notice that in this case a field of moduli is a number field and so is a field of definition of the curves involved in the construction of $X$ too. This yields that a field of definition of the image of $X$ is a number field. As explained in the proof of Theorem 4.1.
Then we consider a linear projection $\pi$ from a generic center of projection defined over $\overline{\mathbb{Q}}$ to $\mathbb{P}^{2}$ Therefore, we get a branch curve, also defined over $\overline{\mathbb{Q}}$ that is irreducible with at most nodes and ordinary cusps as singularities by Theorem 3.2
Using $\tilde{\sigma}$ we get another projection $\pi^{\sigma}\colon X^{\sigma}\longrightarrow\mathbb{P}^{2}$ branched along $B^{\sigma}$. Then by our previous result [LP20] $(\mathbb{P}^{2},B)$ and $(\mathbb{P}^{2},B^{\sigma})$ is a Zariski pair, and by construction they are conjugate, therefore they are an arithmetic Zariski pairs.
Now, consider $<\sigma>\leq{\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}$ the whole cyclic subgroup of order $p$ generated by $\sigma$. Its action on connected components of equisingular strata has orbits only of length $1$ or $p$. By the argument above the orbit of the connected component containing $B$ has at least a second connected component, namely the one containing $B^{\sigma}$. Accordingly the orbit of $B$ under the action of $<\sigma>$ consists of $p$ elements which form an arithmetic Zariski $p$-mulitplet. ∎
Finally we use the article [GJT18] to give more concrete examples of larger multiplets. Their construction involves Beauville surfaces, which are rigid SIP.
@@
$\deg(p_{i})=60p_{i}\left(\frac{p_{i}-1}{6}-1\right)\left(\frac{p_{i}+1}{4}-1\right).$
@@
$(2,3,p-1),\quad(2,4,p+1)\quad\textrm{ and group }{\mathrm{PGL}}(2,p)$
@@
$\lambda(p)=\frac{1}{8}\cdot\varphi\left(p^{2}-1\right)$
where $\varphi$ is the Euler Totien function.
We use the formula given in Lemma 3.10 to get that the degrees of the corresponding branch curves are
$\deg(p)=60p\left(\frac{p-1}{6}-1\right)\left(\frac{p+1}{4}-1\right).$
@@
Assume the contrary that they belong to the same stratum then by our result also the surfaces have isomorphic fundamental group, if so by the weak rigidity theorem they are in the same component contradicting our assumption. ∎

### v2

# A faithful action of Gal($\overline{\mathbb{Q}}/\mathbb{Q}$) on Zariski multiplets

@@ Abstract.
Key words and phrases:
Zariski pairs, absolute Galois group, moduli of surfaces of general type
@@ 1. Introduction
The notion of Zariski pair was introduced by E. Artal in his seminal paper [A94].
Definition 1.1.
A pair of complex reduced plane curves $(B_{1},B_{2})$ is called a Zariski pairs if it satisfies the following conditions:
  1. (1)
    There exist tubular neighborhoods $T(B_{i})$, $i=1,2$, and a homeomorphism $h\colon T(B_{1})\longrightarrow T(B_{2})$ such that $h(B_{1})=B_{2}$.
  2. (2)
    There exists no homeomorphism $f:\mathbb{P}^{2}\longrightarrow\mathbb{P}^{2}$ with $f(B_{1})=B_{2}$ i.e., the pairs $(\mathbb{P}^{2},B_{1})$ and $(\mathbb{P}^{2},B_{2})$ are not homeomorphic.
Analogously $(B_{1},\ldots,B_{k})$ is a Zariski $k$-plet if $(B_{i},B_{j})$ is a Zariski pair for any i $\neq j$.
As noted in [ACT10, Remark 3]) the first condition in Definition 1.1 can be replaced by equality of combinatorial data associated to the $B_{i}$, $i=1,2$, which in case of irreducible curves consist of the degree and the counting function associating to each topological type of singularities the number of corresponding singular points.
This definition put a new focus on an important question for plane complex projective curves:
Do the topological types of the singularities of a plane curve $B$ and the degree determine the homeomorphism type of the pair $(\mathbb{P}^{2},B)$?
@@
In our previous work [LP20] we used the fundamental group of the complement together with the conjugacy class of meridians to the curve, which is well-defined for irreducible curves, in the second step.
@@
In particular we can derive from that set-up the existence of arbitrarily large arithmetic Zariski multiplets of plane curves. Indeed, we are able to make use of specific examples of large Galois orbits of surfaces for the final result.
@@
$\deg(B_{i})=60p_{i}\left(\frac{p_{i}-1}{6}-1\right)\left(\frac{p_{i}+1}{4}-1\right).$
@@
In the final section we first demonstrate that the associated generic $m$-canonical branch curves of surfaces isogenous to a product with different fundamental group form a Zariski pair (see Theorem 3.7) and, using, in particular cases, the functor $F_{\sigma}$, even arithmetic Zariski pairs. Finally, we provide detailed proofs of the main theorems stated in the introduction.
@@ 2. Preliminaries
$F_{\sigma}\colon\underline{\mathbf{\mathbb{C}-var}}$ $\longrightarrow$ $\underline{\mathbf{\mathbb{C}-var}}$
$X$ $\mapsto$ $X^{\sigma}$
$\left(f\colon X\longrightarrow Y\right)$ $\mapsto$ $\left(f^{\sigma}\colon X^{\sigma}\longrightarrow Y^{\sigma}\right)$
@@
and the family $\mathcal{X}^{\sigma}\longrightarrow B^{\sigma}$ connects the surfaces $S_{1}^{\sigma}$ and $S_{2}^{\sigma}$. More precisely, choose the canonical model $X$ of a surface of general type $S$ and apply the field automorphism $\sigma\in{\mathrm{Aut}}(\mathbb{C})$ to a point of the Hilbert scheme corresponding to the $m$-canonical image of $S$ . We obtain a surface $X^{\sigma}$, and whose minimal model is $S^{\sigma}$. Since the Hilbert scheme corresponding to $m$-pluricanonical embedded surfaces is defined over $\mathbb{Z}$, it is invariant under the functor $F_{\sigma}$ so its connected components are defined over $\overline{\mathbb{Q}}$. We conclude that the elements in the kernel of $\varphi\colon{\mathrm{Aut}}(\mathbb{C})\longrightarrow{\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}}$ act trivially on connected components of the moduli space of surfaces which is in bijection with the set of connected components of the Hilbert subscheme just defined. For more details see [BCG15, Proposition 5.7].
We want to do something similar for parameter spaces of plane curves. More precisely, to replace the Hilbert scheme we consider equisingularity strata of plane curves with a fixed combinatorial type. Like the Hilbert scheme, such strata are in general neither irreducible nor connected.
@@
Given a combinatorial type of degree $d$ plane curves as defined in full generality in [ACT10], the corresponding equisingular stratum in the parameter space of degree $d$ plane curve is given by the locus of all curves with that combinatorial type.
Below we will only consider irreducible plane curves with ordinary nodes and cusps with topological type determined by the respective Milnor numbers $\mu=1$ and $\mu=2$. In this situation, the combinatorial datum of $B$ is simply encoded by the triple $(d,n,c)$, where $d=\deg B$ is the degree, $n$ is the number of nodes, and $c$ is the number of cusps of $B$.
The corresponding equisingular stratum is a quasi-projective variety
$\mathcal{S}^{d}_{n,c}\subset\mathbb{P}^{\frac{d(d+3)}{2}}=:\mathbb{P},$
as proved by Wahl in [W74]. In particular, $\mathcal{S}^{d}_{n,c}$ is a locally closed subset of a projective space, and hence a complex algebraic variety endowed with the induced standard analytic topology. Since such spaces are locally path connected, the connected components of $\mathcal{S}^{d}_{n,c}$ coincide with its path-connected components. Consequently, $\mathcal{S}^{d}_{n,c}$ has finitely many (path-)connected components.
Explicitly, we can describe
[…]
We observe that if $(B_{1},B_{2})$ is a Zariski pair, where each $B_{i}$ is given by an equation $\{p_{i}=0\}$, $i=1,2$, then the corresponding points $p_{1}]$ and $p_{2}]$ lie in different path-connected components of $\mathcal{S}^{d}_{n,c}$. The converse is not known in general.
This motivates the following definition where, by abuse of notation, we shall consider a curve $B_{i}$ also as element in $\mathcal{S}^{d}_{n,c}$.
@@
Two curves $B_{1},B_{2}$ in $\mathcal{S}^{d}_{n,c}$ are said to be rigidly isotopic if there exists a continuous path (in the standard analytic topology) $\gamma:[0,1]\to\mathcal{S}^{d}_{n,c}$ such that $\gamma(0)=[p_{1}]$ and $\gamma(1)=[p_{2}]$, where $B_{i}=\{p_{i}=0\}$ for $i=1,2$. Equivalently, $B_{1}$ and $B_{2}$ belong to the same path-connected component of the equisingular stratum $\mathcal{S}^{d}_{n,c}$.
Remark 2.6.
To get this definition into the right perspective let us note the following:
  1. (1)
    In [ACT10], rigid isotopy is defined using smooth paths. We observe that, in this context, the two definitions – using either continuous paths or smooth paths – define the same path components, hence the same equivalence relation.
  2. (2)
    From an algebraic viewpoint, such a path corresponds to a flat projective family of plane curves $\pi:\mathcal{X}\to T$ over a connected complex curve $T$, whose fibers $\mathcal{X}_{t}\subset\mathbb{P}^{2}$ all have degree $d$ and precisely $n$ nodes and $c$ cusps. Thus, rigid isotopy classes coincide with the connected components of the equisingular stratum, both topologically and algebraically.
  3. (3)
    Consequently, the connected components of the equisingular stratum $\mathcal{S}^{d}_{n,c}$ are precisely the rigid isotopy classes of curves.
We now unpack the last two remarks more carefully, since they involve both the analytic and the Zariski topologies.
Let $p_{1}],[p_{2}]\in\mathcal{S}^{d}_{n,c}\subset\mathbb{P}^{\frac{d(d+3)}{2}}$. We define the following equivalence relations reflecting the two topologies:
  1. (1)
    $p_{1}]\sim_{1}[p_{2}]$ if they lie in the same path-connected component of $\mathcal{S}^{d}_{n,c}$ in the standard analytic topology;
  2. (2)
    $p_{1}]\sim_{2}[p_{2}]$ if there exists a finite chain of quasi-projective irreducible curves $T_{1},\dots,T_{m}\subset\mathcal{S}^{d}_{n,c}$ such that $T_{i}\cap T_{i+1}\neq\varnothing$ for all $i$ and $p_{1}],[p_{2}]$ are points on $T_{1}$ resp. $T_{m}$.
    (Equivalently, $\sim_{2}$ is the transitive closure of the relation “lies on a common irreducible quasi-projective curve contained in $\mathcal{S}^{d}_{n,c}$”).
Then these two equivalence relations coincide:
$p_{1}]\sim_{1}[p_{2}]\quad\Longleftrightarrow\quad[p_{1}]\sim_{2}[p_{2}].$
Indeed, the implication from $\sim_{2}$ to $\sim_{1}$ is clear, since each irreducible algebraic curve $T_{i}$ is path connected with the analytic topology.
For the converse, assume $p_{1}]\sim_{1}[p_{2}]$ and let $\gamma:[0,1]\to\mathcal{S}^{d}_{n,c}$ be a path joining them, continuous in the analytic topology of $\mathcal{S}^{d}_{n,c}$. Denote by $S_{1},\dots,S_{\ell}$ the irreducible components of $\mathcal{S}^{d}_{n,c}$ with respect to the Zariski topology. Then for every path $\gamma$ as above, there exist a sequence $i_{1},\dots,i_{k}$ of elements in $\{1,\dots,\ell\}$ and real numbers $0=t_{0}<t_{1}<\dots<t_{k}=1$ such that $\gamma([t_{j-1},t_{j}])\subset S_{i_{j}}$.
By induction on $k\geq 1$ the claim then follows: If $k=1$, both $p_{1}],[p_{2}]$ belong to the same $S_{i_{1}}$ and an irreducible algebraic curve exists on $S_{i_{1}}$ which contains both.
If $k>1$ then by induction hypothesis both $p_{1}],\gamma(t_{k-1})$ belong to a connected union of algebraic curves. Since $\gamma(t_{k-1}),[p_{2}]$ both belong to the same component $S_{i_{k}}$ the inductive claim also follows, thus $p_{1}]\sim_{2}[p_{2}]$.
Moreover, if $p_{1}]\sim_{2}[p_{2}]$ via a connected quasi-projective curve $T\subset\mathcal{S}^{d}_{n,c}$, we can construct the corresponding flat family as follows: consider the universal family $\mathcal{C}\subset\mathbb{P}^{2}\times\mathbb{P}^{\frac{d(d+3)}{2}}$ of irreducible degree-$d$ curves, and let
(2) $\pi:\mathcal{X}:=\mathcal{C}\times_{\mathbb{P}^{\frac{d(d+3)}{2}}}T\longrightarrow T$
be the induced family. Then $\pi$ is a flat projective family of plane curves of degree $d$ with exactly $n$ nodes and $c$ cusps as singularities.
In a similar way as for surfaces we use the diagram (1) in this context, in particular we refer to the map (2) and get:
@@
We want to link more tightly the action of $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}$} on the connected components of moduli space of surfaces with the action on the connected components of equisingular strata of plane curves with only nodes and cusps as singularities. To do so we have to recall the general theory of branched curves, this is the same strategy we used in our previous paper [LP20].
@@
[…]
@@
We are now able to associate to a surface of general type a plane curve with only nodes and cusps as singularities. Among all the surfaces of general type there are some for which it is possible to control accurately the $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}$} action on the connected components of their moduli space. These surfaces are the so called surfaces isogenous to a product (or SIPs for short). They were introduced by Catanese in [Cat00] and the action of $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q)}$} on their moduli space was investigated in [BCG15].
@@
(5) $1\longrightarrow\pi_{1}(C_{1})\times\pi_{1}(C_{2})\longrightarrow\pi_{1}(S)\longrightarrow G\longrightarrow 1.$
@@
is diffeomorphic to $S$. The corresponding moduli space $\mathcal{M}^{top}(S)=\mathcal{M}^{\it diff}(S)$ of surfaces (orientedly) homeomorphic (resp. diffeomorphic) to $S$ is either irreducible and connected or consists of two irreducible connected components exchanged by complex conjugation.
@@ 4. Proof of the main theorems
By Theorem 3.7 $(\mathbb{P}^{2},B)$, $(\mathbb{P}^{2},B^{\sigma})$ are Zariski pair. We only have to prove that the pair is arithmetic. By the proof of [BCG15, Theorem 1.3] the field of moduli is a number field. Also a field of definition is a number field, though a bigger one in general, therefore we know that the $m$-canonical model $X$ of $S$ and $X^{\sigma}$ for $S^{\sigma}$ in $\mathbb{P}^{P_{m}-1}$ is defined over $\overline{\mathbb{Q}}$. By [CiF11, Theorem 1.1] we can map $S$ to $\mathbb{P}^{2}$ by a generic projection $\pi$ with branch curve $B$ that is irreducible with at most nodes and ordinary cusps as singularities and again defined over a number field. In particular, this allows us to apply the functor $F_{\sigma}$ defined in 2.2 and to obtain the following diagram.
${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${X^{\sigma}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{\sigma}}$$\scriptstyle{F_{\sigma}}$${B\subset\mathbb{P}^{2}}$${B^{\sigma}\subset\mathbb{P}^{2}}$
@@
We are now in the position to prove Theorems 1.3 and 1.4, that we recall for reader’s convenience.
@@
Finally we use the article [GJT18] to give more concrete examples of larger multiplets. Their construction involves Beauville surfaces, which are rigid SIP and their result is stated in terms of the Euler totient function $\varphi$:
@@
$\deg(B_{i})=60p_{i}\left(\frac{p_{i}-1}{6}-1\right)\left(\frac{p_{i}+1}{4}-1\right).$
@@
$(2,3,p-1),\quad(2,4,p+1)\quad\textrm{ and group}\quad{\mathrm{PGL}}(2,p)$
@@
$\lambda(p)=\frac{1}{8}\cdot\varphi\left(p^{2}-1\right).$
We use the formula given in Lemma 3.10 to get that the degrees of the corresponding branch curves $B_{p}$ are
$\deg(B_{p})=60p\left(\frac{p-1}{6}-1\right)\left(\frac{p+1}{4}-1\right).$
@@
Assume the contrary that they belong to the same stratum then by our result [LP20, Theorem 3.12] also the surfaces have isomorphic fundamental group, if so by the weak rigidity theorem they are in the same component contradicting our assumption. ∎